OpenStudy (anonymous):

Arnold is about to go on a 500-mile car trip. His mechanic recommends that he buy a special highway engine oil that will save him 50 cents in gas for every 25 miles of the trip. This new oil, however, will cost $20. Is it worthwhile for Arnold to buy the oil if he has a coupon for $4 dollars off the price?

OpenStudy (anonymous):

total money saved = 500/25 x 50 = 1000 cent =$10 price of oil = 20-4 = $16 therefore, he has to pay $6 more. Is it not worthwhile